OpenStudy (anonymous):

2/h-7=12/5h-2h+3 HELP !!! PLZ

OpenStudy (anonymous):

HELLP !!! PLZ

OpenStudy (anonymous):

hello

OpenStudy (yttrium):

Is that what you mean? \[\frac{ 2 }{ h-7 } = \frac{ 12 }{ 5h } - 2h+3\]

OpenStudy (anonymous):

no its \[\frac{ 2 }{ 5 }h - 7 = \frac{ 12 }{ 5 }h - 2h + 3\]

OpenStudy (yttrium):

You can multiply the whole equation by 5 so that it will be cancelled. After that is the time to solve for h. :) You follow?

OpenStudy (anonymous):

no srry can u do it step by step plz :D

OpenStudy (yttrium):

Let me start with what I've said It will appear like this \[5[\frac{ 2h }{ 5 }-7 = \frac{ 12h }{ 5 } - 2h +3]\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

wait

OpenStudy (yttrium):

Do you have idea now?

OpenStudy (anonymous):

its not \[\frac{ 2h }{ 5 } its \frac{ 2 }{ 5 }h\]

OpenStudy (yttrium):

Don't worry, they are just the same. :) Do you know what to do next?

OpenStudy (anonymous):

oh :D

OpenStudy (yttrium):

Then distributing the 5 will cancel the denominator. And this will happen next 2h - 7(5) = 12h - 2h (5) + 3(5) just solve for h and that's it :)

OpenStudy (anonymous):

can u slove it i dont get

OpenStudy (yttrium):

2h - 35 = 12h - 10h +15 So, do you know how to continue it?

OpenStudy (anonymous):

subtract 15 from bot sides then what

OpenStudy (yttrium):

Wait, do you know something about axioms? Or transpositions?

OpenStudy (anonymous):

what ?

OpenStudy (yttrium):

nevermind. i think it's better if you isolate first the variable against the constants.

OpenStudy (yttrium):

Do you know it?